# Federated Instruction Tuning of LLMs with Domain Coverage Augmentation

### v1

@@ Abstract
Federated Domain-specific Instruction Tuning (FedDIT) leverages a few cross-client private data and server-side public data for instruction augmentation, enhancing model performance in specific domains. While the factors affecting FedDIT remain unclear and existing instruction augmentation methods mainly focus on the centralized setting without considering the distributed environment. Firstly, our experiments show that cross-client domain coverage, rather than data heterogeneity, drives model performance in FedDIT. Thus, we propose FedDCA, which maximizes domain coverage through greedy client center selection and retrieval-based augmentation. To reduce client-side computation, FedDCA uses heterogeneous encoders with server-side feature alignment. Extensive experiments across four domains (code, medical, financial, and mathematical) validate the effectiveness of both methods. Additionally, we explore the privacy protection against memory extraction attacks with various amounts of public data and results show that there is no significant correlation between the amount of public data and the privacy-preserving capability. However, as the fine-tuning round increases, the risk of privacy leakage reduces or converges.
@@ 1 Introduction
Table 1: Performance(%) of different augmentation settings on each domain, which are conducted via FedAvg protocol with 10 clients. Specifically, the zero-shot directly inferences without fine-tuning, while base data utilizes only the client’s local data for FedDIT. Also, we compare FedDCA with other two augmentation strategies: random sampling and direct retrieval (described in Appendix A.3), respectively.
@@
Recently, federated instruction tuning (FedIT) has gained attention as a novel approach that leverages the principles of federated learning (FL) to enable collaborative training of large language models (LLM) in a distributed environment without the need to share private data (McMahan et al. (2017); Ye et al. (2024b)). This approach involves exchanging model parameters among distributed data holders, striking a delicate balance between preserving privacy and optimizing model efficiency. Although many FedIT frameworks have been established(Ye et al. (2024b); Kuang et al. (2023); Zhang et al. (2024a)), they have not deeply explored the practical issues that FedDIT might encounter in real-world scenarios. For example, FedIT usually requires sufficient instruction data for fine-tuning, which is difficult to meet in domain-specific fine-tuning scenarios (Zhang et al. (2024d)).
@@
In this study, we explore Federated Domain-specific Instruction Tuning (FedDIT), a novel approach within the FL paradigm that focuses on enhancing the performance of LLMs in specific domains. Unlike the general FedIT, which seeks to improve model performance across a variety of tasks without considering local data shortage, FedDIT faces the challenge of clients having only a few local domain-specific data. To overcome this, FedDIT leverages a server-hosted public dataset that spans multiple domains to enrich the local data through the specific instruction augmentation strategy. This strategic enrichment is essential for achieving effective instruction tuning and needs to be carefully designed to prevent performance degradation. Except for the code domain, which is more about following a certain paradigm. As shown in Table 1, when clients attempt to do FedDIT only based on their local data, although some high-quality in-domain local data may be available, the limited scale restricts effective instruction tuning. Consequently, this sometimes leads to performance degradation, as indicated by the underlined values. Overall, the goal of FedDIT is to develop a domain-specific LLM that utilizes collaborative training and instruction augmentation with public data while preserving client privacy, ensuring that the model is adept at handling tasks within its targeted domain.
@@
Additionally, the factors affecting FedDIT are still unclear. What’s worse, augmenting instructions may make the results even more unpredictable and confused about how to improve. Shepherd(Zhang et al. (2024a)) approaches this problem from the perspective of heterogeneity, constructing heterogeneity based on the topic of general instruction datasets. The study demonstrates that, for general FedIT, unlike the consensus of traditional FL, heterogeneity has a positive effect. By aggregating diverse instructions from clients, the diversity increases, thereby enhancing the model’s adaptability to various tasks. However, it just scratches the surface and does not explore issues in FedDIT.
@@
Going one step further, we conduct experiments to unveil a new key finding (Appendix A.2): there is no linear correlation between the degree of non-independent and identically distributed (non-iid) and the performance of LLM for FedDIT. Inspired by Explore-Instruct(Wan et al. (2023)), we first define the domain coverage metric and investigate the effect of domain coverage. Results show that the cross-client domain coverage impacts the model’s performance significantly.
@@
To maximize the cross-client domain coverage without compromising client data privacy, based on INSTA(Lee et al. (2024)), we propose a novel FedDIT algorithm, called FedDCA (Federated Instruction Tuning of LLMs with Domain Coverage Augmentation), which performs greedy client center selection and instruction augmentation via dense retrieval on the server side. The core idea of FedDCA lies in greedy client center selection, which chooses client centers that enhance the diversity and coverage of the domain-specific characteristics within the public dataset. By concentrating on maximizing domain coverage at each step, the algorithm efficiently builds an augmented train set that boosts the model’s learning and generalization capabilities, resulting in superior performance on domain-specific tasks. Furthermore, to reduce the computation overhead on the client side, we propose FedDCA with heterogeneous encoders. For feature alignment, we train a projector on the server side using public data through contrastive learning.
@@
We demonstrate the effectiveness of FedDCA by conducting extensive experiments across four domains: code, medical, financial, and mathematical, and comparing against multiple baselines, which can be categorized into two types: unaugmented and augmented methods. For the unaugmented setting, we compare our method with FedIT, which composes of four orthodox FL methods, including FedAvg(McMahan et al. (2017)), FedProx(Li et al. (2020)), SCAFFOLD(pmlr-v119-karimireddy20a), and FedAvgM(Hsu et al. (2019)). For the augmented setting, we compare our method with random sampling, direct retrieval, LESS(Xia et al. (2024)), and Self-Instruction(Wang et al. (2023c)). Additionally, we report FedDCA’s performance with different FL strategies. Furthermore, we compare the computational efficiency of FedDCA and FedDCA on the client side. To illustrate the impact of different retrieval volumes on client data privacy, we conduct experiments against memory extraction attacks. Results indicate that as the number of training rounds increases, the method only based on the client’s local data memorizes more privacy information while the privacy leakage risk gradually decreases or converges for the augmented setting.
@@
  • We unveil a key finding: For FedDIT, there is no linear correlation between the degree of heterogeneity and the model’s performance. Instead, cross-client domain coverage impacts the model’s performance significantly, which is presented in Appendix A.2.
  • We propose a novel FedDIT algorithm (Section 4) called FedDCA to maximize the cross-client domain coverage through greedy client center selection and then perform retrieval-based instruction augmentation on the server. To further reduce the computational overhead on the client side, we propose FedDCA with a heterogeneous encoder setting. For feature alignment, we train a projector on the server side.
  • With extensive experiments (Section 5), we show the effectiveness of FedDCA and FedDCA. Furthermore, we explore privacy preservation against memory extraction attacks and conduct experiments based on various public data amounts. Results indicate that the privacy-preserving capability is not significantly correlated with the amount of public data, but the risk of privacy leakage decreases or converges as the amount of fine-tuning rounds increases.
@@ 2 Preliminaries
Federated domain-specific instruction tuning (FedDIT) aims to leverage cross-client private domain-specific instruction data and utilize the multi-domain public data on the server to achieve instruction augmentation, collaboratively enhancing the model’s performance in specific domains. Suppose $N$ distributed clients with their local private data $D_{k}^{l}$. The server has a public dataset $D^{p}$ consisting of multiple domains and is responsible for data augmentation based on different strategies and aggregation model parameters from clients. The training process follows standard FL protocol(McMahan et al. (2017)). For computational efficiency, we adopt LoRA(Hu et al. (2021)) as the fine-tuning method, which involves tuning additional parameters $\Delta\phi$ while keeping the pre-trained LLM’s parameters $\phi$ frozen. In the initial training stage, the server dispatches the pre-trained LLM parameters $\phi$ and the initialized LoRA parameters $\Delta\phi$ to each client for training $\mathcal{R}$ rounds. In the $t$-th round, the server sends the aggregated LoRA parameters $\Delta\phi^{t}$ to clients. Clients use $\Delta\phi^{t}$ to update their local LoRA parameters $\Delta\phi_{k}^{t}$ and perform instruction tuning based on augmented instruction datasets, then send the updated LoRA parameters $\Delta\phi_{k}^{t+1}$ back to the server. The server receives the LoRA parameters $\{\Delta\phi_{k}^{t+1}\mid k=1,\dots,N\}$ from each client and aggregates them to obtain the global LoRA parameters $\Delta\phi^{t+1}$ for the next round.
@@ 3 Problem Formulation
For better understanding, we list the frequently used notation in Appendix A.11. The objective of FedDIT is to enhance the domain-specific performance of LLMs through FL without sharing private data(Ye et al. (2024b); Zhang et al. (2024d)). Under the FL framework, suppose we have $N$ clients, where each client $c_{k}$ has a local dataset $D_{k}^{l}$ with its size $N_{k}^{l}$, and an augmented dataset $D_{k}^{g}$ from the server public dataset $D^{p}$, respectively. Due to constraints such as memory, computational overhead, and maximum tolerated training time, client $c_{k}$ can accept at most $N_{k}^{p}$ public instructions. Denote the augmentation strategy as $\Lambda$, through which the server performs instruction augmentation on the public dataset. If $\Lambda$ is null, it denotes clients only perform FedDIT based only on their local private data, which may cause performance degradation. Otherwise, $\Lambda$ could be categorized into two types: 1) Focusing only on its own local data distribution. 2) Taking into account the cross-client data distribution. In conclusion, the global objective of FedDIT is defined as follows:
@@
where the first term and second term denote the accumulated fine-tuning loss computed on the client $c_{k}$’s local instructions $D_{k}^{l}$ and the augmented public instructions $D_{k}^{g}$ from the server respectively. $p_{k}$ is the weight of the $k$-th client, which is determined by the ratio of $k$-th client’s data size to all clients’ total data size. $\alpha_{k}$ is the ratio of the public data amount of its augmented instruction dataset $D_{k}$, which is computed as $\frac{N_{k}^{p}}{N_{k}^{l}+N_{k}^{p}}$. Eq.1 minimizes the summed empirical loss across clients’ augmented instructions to pursue the in-domain utility of the obtained global model. Denote $P$ as the model parameters and $\mathcal{D}$ as a specific dataset, then the client $c_{k}$’s empirical loss $F_{k}(\phi,\Delta\phi_{k};\mathcal{D})$ base on $\mathcal{D}$ is calculated as:
@@ 4 Method
We propose Federated Instruction Tuning of LLMs with Domain Coverage Augmentation (FedDCA), which enhances domain coverage through greedy client center selection and performs retrieval-based instruction augmentation to obtain a LLM that performs well on domain-specific tasks. We first give the overview FedDCA (Section 4.1) and define the domain coverage metric (Section 4.2), then show details of the greedy client center selection algorithm (Section 4.3) and heterogeneous encoder method FedDCA (Section 4.4).
4.1 Overview
Figure 1: Overview of FedDCA, which consists of three stages: 1) The client $c_{k}$ performs local instructions clustering and sends cluster centers $\mathcal{C}_{k}$ to the server. 2) The server first does greedy client center selection to maximize domain coverage and performs client-center-based domain data retrieval, then sends the augmented instructions $D_{k}^{g}$ to the client $c_{k}$. 3) Clients fine-tune the LLM collaboratively and the LoRA parameters $\Delta\phi$ are exchanged between clients and the server.
As shown in Figure 1, FedDCA consists of three stages, in the first stage the client $c_{k}$ clusters its local instructions and sends the cluster centers $\mathcal{C}_{k}$ to the server. Then the server performs greedy client center selection to obtain client centers that maximize the domain coverage, then performs client-center-based retrieval for domain-specific instruction augmentation. Additionally, the third stage follows the standard FedIT procedure(Ye et al. (2024b)), in which the server sends the frozen pre-trained LLM parameters $\phi$ and the initialized LoRA parameters $\Delta\phi$ to clients and then $\Delta\phi_{k}$ is exchanged between the client and the server.
The novel design of FedDCA lies in the second stage before training. As shown in Algorithm 1, the core idea is to use a greedy client center selection algorithm to obtain client centers that maximize cross-client domain coverage, and then use these client centers to retrieve instructions on the public dataset. To further reduce the computational overhead of encoding on clients, we propose a heterogeneous encoder method called FedDCA. Specifically, the client uses a small encoder locally, while the server uses a larger encoder for accurate semantics. To align features with different dimensions, we introduce a projector $w_{p}$ on the server side, which is obtained through contrastive learning.
@@ 4.1 Overview
4.2 Domain Coverage
As domain coverage directly affects the in-domain performance of the LLM (Appendix A.2), FedDCA aims to maximize the domain coverage of the cross-client data distribution $\cup_{i=1}^{N}D_{i}$ respect to the in-domain data distribution $D^{d}$, which is defined in Eq. 5. However, as public data is unreachable for clients and clients can not send the local data to the server for privacy, directly finding a cross-client dataset that maximizes the domain coverage in Eq. 5 is unrealistic. To find a proper client center set $\mathcal{P}$ that attempts to maximize the domain coverage $d(D^{d},P)$ and use $\mathcal{P}$ to perform instruction retrieval on the public data is an approximation problem.
Suppose there exists a metric space $\mathcal{X}$, a set of cross-client local instruction embeddings $\mathcal{E}\in\mathcal{X}$, and a set of cluster centers $\mathcal{P}\in\mathcal{X}$. Each client $c_{k}$ has maximum $\xi$ clusters obtained by k-means algorithm(Vardakas & Likas (2023)). As in the federated setting, communication cost is always considered. Thus, we need to find a $\mathcal{P}$ that maximizes the domain coverage while considering the communication cost, we formulate the optimization problem as follows:

### v3

@@ Abstract
Federated Domain-specific Instruction Tuning (FedDIT) utilizes limited cross-client private data together with server-side public data for instruction augmentation, ultimately boosting model performance within specific domains. To date, the factors affecting FedDIT remain unclear, and existing instruction augmentation methods primarily focus on the centralized setting without considering distributed environments. Our experiments reveal that the cross-client domain coverage, rather than data heterogeneity, drives model performance in FedDIT. In response, we propose FedDCA, which optimizes domain coverage through greedy client center selection and retrieval-based augmentation. For client-side computational efficiency and system scalability, FedDCA, the variant of FedDCA, utilizes heterogeneous encoders with server-side feature alignment. Extensive experiments across four distinct domains (code, medical, financial, and mathematical) substantiate the effectiveness of both methods. Additionally, we investigate privacy preservation against memory extraction attacks utilizing various amounts of public data. Results show that there is no significant correlation between the volume of public data and the privacy-preserving capability. However, as the fine-tuning rounds increase, the risk of privacy leakage reduces or converges.
@@ 1 Introduction
Table 1: Performance(%) of different augmentation settings on each domain, conducted via FedAvg protocol with 10 clients. Specifically, the zero-shot directly inferences without any fine-tuning, while the base data method utilizes only the client’s local data for FedDIT. Additionally, we compare FedDCA with other two augmentation strategies: random sampling and direct retrieval (described in Appendix A.3), respectively.
@@
Recently, federated instruction tuning (FedIT) has gained attention as a novel approach that leverages the principles of federated learning (FL) to facilitate collaborative training of large language models (LLM) in distributed environments while maintaining the confidentiality of private data (McMahan et al., 2017; Ye et al., 2024b). This methodology allows for the exchange of model parameters among distributed data holders, thereby achieving a careful balance between privacy preservation and efficient model optimization. Despite the establishment of various FedIT frameworks (Ye et al., 2024b; Kuang et al., 2023; Zhang et al., 2024a), existing literature has not adequately addressed the practical challenges that Federated Domain-specific Instruction Tuning (FedDIT) may encounter in real-world applications. For instance, FedIT generally necessitates a sufficient amount of instruction data for fine-tuning, which is often a shortage in domain-specific fine-tuning contexts (Zhang et al., 2024d).
@@
In this study, we investigate FedDIT, a novel approach within the FL paradigm aimed at boosting the performance of LLMs in specific domains. Unlike general FedIT, which seeks to enhance model effectiveness across diverse tasks without accounting for local data shortage, FedDIT encounters the unique challenge of clients possessing only a limited quantity of local domain-specific data. To overcome this, FedDIT leverages a server-hosted public dataset that spans multiple domains to enrich the local data through a specific instruction augmentation strategy. This strategic enrichment is crucial for achieving effective instruction tuning and needs to be meticulously designed to avoid performance degradation. Except for the code domain, which primarily adheres to a standardized paradigm, our results reveal that when clients rely solely on their local data for FedDIT, even the presence of high-quality in-domain local data can be insufficient due to limited scale, leading to a decline in performance, as reflected in the underlined values in Table 1. In summary, the goal of FedDIT is to develop a domain-specific LLM that employs collaborative training and instruction augmentation with public data while safeguarding client privacy, thereby ensuring the model’s proficiency in executing tasks pertinent to its designated domain.
@@
Additionally, the factors affecting FedDIT are still unclear. Compounding this uncertainty, introducing augmented instructions may further complicate results, making it difficult to ascertain effective improvement strategies. Shepherd (Zhang et al., 2024a) approaches this problem from the perspective of heterogeneity, constructing heterogeneity based on the topic of general instruction datasets. It demonstrates that, unlike the consensus of traditional FL, for general FedIT, heterogeneity has a positive effect. By aggregating diverse instructions from clients, the diversity increases, thereby enhancing the model’s adaptability to various tasks. However, it just scratches the surface and does not explore issues in FedDIT.
@@
Going one step further, we conduct experiments to unveil a significant finding (Appendix A.2): there is no linear correlation between the degree of non-independent and identically distributed (non-iid) and LLM’s performance in the context of FedDIT. Inspired by Explore-Instruct (Wan et al., 2023), which shows the potential of domain coverage in domain-specific instruction tuning. The cross-client domain coverage metric is initially defined, followed by an investigation into its impact on FedDIT. Results demonstrate that domain coverage significantly influences model performance in the corresponding domain.
@@
To maximize the cross-client domain coverage without compromising client data privacy, we propose a novel FedDIT algorithm, Federated Instruction Tuning of LLMs with Domain Coverage Augmentation, termed FedDCA. This algorithm employs a greedy client center selection process and implements instruction augmentation through dense retrieval on the server side. The fundamental idea of FedDCA is to select client centers to expand the diversity and coverage of augmented instruction datasets within a specific domain. By strategically optimizing domain coverage at each step, FedDCA efficiently constructs the augmented train set that enhances both the learning and generalization capabilities of the model, leading to superior performance on domain-specific tasks. Furthermore, to mitigate computational overhead on the client side and enhance the system scalability, we propose FedDCA, which employs heterogeneous encoders of different sizes and capacities. To achieve feature alignment, we train a projector on the server side using public data and employ contrastive learning techniques.
@@
We demonstrate the effectiveness of FedDCA through comprehensive experiments conducted across four domains: code, medical, financial, and mathematical. These are compared against a range of baselines, which can be categorized into unaugmented and augmented methods. In the unaugmented setting, our method is compared with FedIT, which includes four orthodox FL techniques: FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020), SCAFFOLD (pmlr-v119-karimireddy20a), and FedAvgM (Hsu et al., 2019). In the augmented setting, we compare FedDCA against methods such as random sampling, direct retrieval, LESS (Xia et al., 2024), and Self-Instruction (Wang et al., 2023c). Additionally, we present the performance outcomes of FedDCA when applied under various FL strategies. We also compare the computational efficiency on the client side between FedDCA and its variant, FedDCA. For privacy analysis, experiments against memory extraction attacks are conducted to evaluate how different quantities of retrieved public data affect the privacy of client local data. Results indicate that while reliance on local data increases memorization of sensitive information, the risk of privacy leakage diminishes or converges in the augmented setting as the training rounds progress.
@@
  • We reveal a critical finding: in the context of FedDIT, there exists no linear correlation between the degree of heterogeneity and model performance. Instead, cross-client domain coverage substantially impacts LLM’s effectiveness, as elaborated in Appendix A.2.
  • We propose a novel FedDIT algorithm (Section 4), termed FedDCA, aimed at maximizing cross-client domain coverage through greedy client center selection followed by retrieval-based instruction augmentation executed on the server. Additionally, we introduce FedDCA to further lessen the client-side computational overhead while enhancing the system scalability. This variant utilizes a heterogeneous encoder structure, paired with a projector on the server side for feature alignment.
  • Through extensive experiments (Section 5), we demonstrate the effectiveness of FedDCA and FedDCA. We also investigate privacy preservation against memory extraction attacks, conducting experiments based on various amounts of public data. Results suggest that the capacity for privacy preservation does not correlate significantly with the quantity of public data. In contrast, the risk of privacy leakage tends to decrease or converge as the fine-tuning rounds increase.
@@ 2 Preliminaries
FedDIT aims to leverage cross-client private domain-specific instruction data and utilize the multi-domain public data on the server to achieve instruction augmentation, collaboratively enhancing the model’s performance in specific domains. Consider $N$ distributed clients, each with local private data $D_{k}^{l}$. The server maintains a public dataset $D^{p}$ that encompasses multiple domains and is responsible for implementing data augmentation strategies and aggregating model parameters received from clients. The training process follows the standard FL protocol (McMahan et al., 2017). For computational efficiency, we adopt Low-Rank Adaption (LoRA) (Hu et al., 2021) as the fine-tuning method, which involves tuning additional parameters $\Delta\phi$ while keeping the pre-trained LLM’s parameters $\phi$ frozen. In the initial training phase, the server dispatches $\phi$ and $\Delta\phi$ to each client for $\mathcal{R}$ training rounds. In the $t$-th round, the server sends the aggregated $\Delta\phi^{t}$ to clients. Clients use $\Delta\phi^{t}$ to update their local LoRA parameters $\Delta\phi_{k}^{t}$ and conduct instruction tuning based on augmented instruction datasets $D_{k}$. Subsequently, the clients return $\Delta\phi_{k}^{t+1}$ to the server. The server then aggregates $\{\Delta\phi_{k}^{t+1}\mid k=1,\dots,N\}$ to obtain $\Delta\phi^{t+1}$ for the next round.
@@ 3 Problem Formulation
For better understanding, we list the frequently used notation in Appendix A.11. The objective of FedDIT is to enhance the domain-specific performance of LLMs through FL without sharing private data (Ye et al., 2024b; Zhang et al., 2024d). Under the FL framework, suppose we have $N$ clients, where each client $c_{k}$ has a local dataset $D_{k}^{l}$ with its size $N_{k}^{l}$, and an augmented dataset $D_{k}^{g}$ from the server public dataset $D^{p}$, respectively. Due to constraints such as memory, computational overhead, and maximum tolerated training time, client $c_{k}$ can accept at most $N_{k}^{p}$ public instructions. Denote the augmentation strategy as $\Lambda$, through which the server performs instruction augmentation on the public dataset. If $\Lambda$ is null, it indicates that clients conduct FedDIT solely based on their local private data, which may lead to performance degradation. Conversely, $\Lambda$ may be classified into two categories: (1) focusing exclusively on the client’s own local data distribution or (2) considering the cross-client data distribution. In conclusion, the global objective of FedDIT is defined as follows:
@@
where the first term and second term denote the accumulated fine-tuning loss computed on the client $c_{k}$’s local instructions $D_{k}^{l}$ and the augmented public instructions $D_{k}^{g}$ from the server, respectively. $p_{k}$ is the weight of the $k$-th client, which is determined by the ratio of $k$-th client’s data size to all clients’ total data size. $\alpha_{k}$ is the ratio of the public data amount of its augmented instruction dataset $D_{k}$, which is computed as $\frac{N_{k}^{p}}{N_{k}^{l}+N_{k}^{p}}$. Eq.1 minimizes the summed empirical loss across clients’ augmented instructions to pursue the in-domain utility of the obtained global model. Denote $P$ as the model parameters and $\mathcal{D}$ as a specific dataset, then the client $c_{k}$’s empirical loss $F_{k}(\phi,\Delta\phi_{k};\mathcal{D})$ base on $\mathcal{D}$ is calculated as:
@@ 4 Method
We propose Federated Instruction Tuning of LLMs with Domain Coverage Augmentation (FedDCA), which enhances domain coverage to obtain a LLM that performs well on domain-specific tasks. FedDCA follows the standard FedIT protocol(Ye et al., 2024b) to perform federated instruction tuning. The novel design of FedDCA lies in the phase before training, which consists of two key modules: greedy client center selection and domain data retrieval, which are elaborated in the following; also see details in Algorithm 1 and Figure 1. For computational efficiency and system scalability, we introduce a variant of FedDCA with heterogeneous encoder setting, named FedDCA. We first formulate the optimization problem to solve for FedDCA.
Figure 1: Overview of FedDCA, which consists of three stages: 1) The client $c_{k}$ performs local instructions clustering and sends cluster centers $\mathcal{C}_{k}$ to the server. 2) The server first does greedy client center selection to maximize domain coverage and performs client-center-based domain data retrieval, then sends the augmented instructions $D_{k}^{g}$ to the client $c_{k}$. 3) Clients fine-tune the LLM collaboratively, and the LoRA parameters $\Delta\phi$ are exchanged between clients and the server.
@@
4.1 Optimization Problem
As domain coverage directly affects the in-domain performance of the LLM (Appendix A.2), FedDCA aims to maximize the domain coverage of the cross-client augmented data $\cup_{i=1}^{N}D_{i}$ respect to the in-domain data distribution $D^{d}$, which is defined in Eq. 5. However, as clients can not send the local data to the server for privacy, directly finding a cross-client dataset that maximizes the domain coverage in Eq. 5 is unrealistic. To find a proper client center set $\mathcal{P}$ that maximizes domain coverage and uses it to perform instruction retrieval on the public data is an approximation problem.
Suppose there exists a metric space $\mathcal{X}$, a set of cross-client local instruction embeddings $\mathcal{E}\in\mathcal{X}$, and a set of cluster centers $\mathcal{P}\in\mathcal{X}$. Each client $c_{k}$ has maximum $\xi$ clusters obtained by k-means algorithm (Vardakas & Likas, 2023). In the federated setting, communication cost is always a critical factor. Consequently, we formulate the optimization problem to include communication costs as follows: